OpenStudy (pastorswife3):

What is the solution for the following system of linear equations? x + 6y + -4 and 3x - 4y = 10. Show the work so I will understand it

OpenStudy (anonymous):

x+6y=-4 x=-4-6y put value of x in other equation 3*(-4-6y)-4y=10 -12-18y-4y=10 -22y=22 y= -1

OpenStudy (anonymous):

put y =-1 in x=-4-6y=-4-6*(-1)=-4+6=2 x=2 y=-1

OpenStudy (anonymous):

see the solution in the attached pdf file